# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

### v2

@@ E.3 Result Tables
Downstream Generator Predicted / Measured
@@
Downstream Generator Predicted / Measured
@@
Gradient Boosting AIM - Log Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
@@ E.4 Estimating Model and Synthetic Data Variances
In this section, we estimate the MV and SDV terms from the decomposition in Theorem 2.2. We first generate 32 synthetic datasets that are 5 times larger than the real dataset, and split each synthetic datasets into 5 equally-sized subsets. This is equivalent to sampling 32 $\theta_{i}$ values, and for each $\theta_{i}$, generating 5 synthetic datasets i.i.d.  We then train the downstream predictor on each synthetic dataset, and store the predictions for all test points.

### v3

@@ E.3 Result Tables
Downstream Generator
@@
Downstream Generator
@@
Grad. Boosting AIM - Log Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
@@ E.4 Estimating Model and Synthetic Data Variances
In this section, we estimate the MV and SDV terms from the decomposition in Theorem 2.2. We first generate 32 synthetic datasets that are 5 times larger than the real dataset, and split each synthetic datasets into 5 equally-sized subsets. This is equivalent to training 32 generators, with parameters $\theta_{i}$, and for each generator, generating 5 synthetic datasets i.i.d.  We then train the downstream predictor on each synthetic dataset, and store the predictions for all test points.
@@
E.5 Comparison with Generating A Single Large Synthetic Dataset
In this section, we examine an alternative to multiple synthetic datasets: generating a single, large synthetic dataset. van Breugel et al. (2023a) found this lead to poor model evaluation and selection, with a single synthetic dataset leading to overestimating the performance of complex models. They did not directly compare the effect on accuracy, which is what we will do in this section.
We use the synthetic datasets from the variance estimation experiment [^14] We use only one of the 32 synthetic datasets per repeat. of Section E.4 that are 5 times larger than the real dataset. We also take smaller subsets of the whole synthetic dataset to examine other synthetic dataset sizes. We train the same models as in the other experiments, but only on regression datasets generated by synthpop (proper).
We compare the results from the single large synthetic dataset with the ensemble of multiple synthetic dataset with an equal number of total datapoints in Figure S12. We see that the ensemble of multiple synthetic datasets is equal or better in all cases in terms of MSE.
(a) Abalone
(b) ACS 2018
(c) California Housing
(d) Insurance
Figure S12: Comparison of the ensemble of multiple synthetic datasets with one large synthetic dataset with an equal number of synthetic datapoints on the regression datasets with synthpop. We see that multiple synthetic datasets are always equal or better. Size Mul. is the relative total number of synthetic data points to the real data: for multiple synthetic datasets, it is $m$, and for a single large synthetic dataset it is $n_{Syn}/n_{Real}$.